# Convexity on manifolds without focal points and spectrum of the Laplacian

### v1

@@ Abstract.
Convexity on complete Riemannian manifolds has typically been studied under constraints on the sign of sectional curvatures. This is especially so in the study of the effect of convexity on the spectrum of the Laplacian. In this article, we use the language of Jacobi tensors to construct a strictly convex function with bounded gradient on complete simply connected manifolds without focal points having negative Ricci curvature. From this we obtain a lower bound for the growth of the volume of geodesic balls, if the manifold is also assumed to be Kähler. Our primary applications concern the spectrum of the Laplacian. We use the strictly convex function which we construct to show that the spectrum of the Laplacian is absolutely continuous on a class of asymptotically harmonic manifolds. In fact, we completely determine the spectrum of the Laplacian on simply connected noncompact nonflat harmonic manifolds without focal points. For asymptotically harmonic manifolds we also show that equality holds in Cheeger’s inequality for the bottom of the spectrum, where Cheeger’s constant is equal to the constant mean curvature of the horospheres. We show that the well-known description of the spectrum of the Laplacian on Riemannian symmetric spaces of noncompact type and arbitrary rank can also be recovered using these geometric methods.
@@ 1. Introduction
Convex functions have long played a fundamental role in Riemannian geometry. By the convexity of a continuous real-valued function on a Riemannian manifold we mean that it is convex in the usual sense along any geodesic. The mere existence or nonexistence of convex functions have geometric and topological implications. Nonconstant convex functions on complete Riemannian manifolds exist only on those with infinite volume [2, Proposition 2.2]. In particular, they don’t exist on compact Riemannian manifolds. There are profound structural results arising from the interaction of curvature and convexity. Under appropriate curvature conditions there are natural geometrically defined convex functions on a Riemannian manifold. On complete simply connected Riemannian manifolds of nonpositive sectional curvature, called Hadamard manifolds, the square of the distance function from a fixed but arbitrary point and the Busemann functions are examples of convex functions which play a crucial role in the geometry of these manifolds. While the former is a smooth convex function the latter is merely a $C^{2}$-convex function [12, §1.6.6]. In the nonnegative sectional curvature situation for instance, a key idea in the fundamental work [8] is the construction of suitable continuous convex functions related to Busemann functions which have appropriate additional properties.
@@
A function which is at least $C^{2}$ is (strictly) convex if its hessian is positive (definite) semi-definite everywhere. If a Riemannian manifold admits a strictly convex function, then it has no compact immersed minimal submanifolds [20, §9.2 C]. In particular, it has no closed geodesics. Such a manifold also allows for a generalization of Cartan’s theorem [29] when this strictly convex function has a minimum: every compact subgroup of the isometry group has a common fixed point. The existence of strict convex functions also has consequences for the $\mathrm{L^{2}}$-spectrum of the Laplacian. The study of the spectrum of the Laplacian on complete Riemannian manifolds has a long, rich history and there are very many results, often with various additional constraints placed on the structure of the manifold and on sectional curvature. For a succint recent overview of some of these aspects, we refer to [9]. We content ourselves by mentioning a few results which use convexity in an essential manner to deduce information on the spectrum to put our work in context, with no pretensions to completeness.
The absolutely continuous part of the spectrum of the Laplacian on complete Riemannian manifolds supporting a sufficiently regular strictly convex function is nontrivial [28, Theorem 2] under certain boundedness conditions on the gradient and the Laplacian of the said function. On complete simply connected manifolds with nonpositive sectional curvature, [28, Theorem 1] exhibited a strictly convex function related to the square of the distance function satisfying the estimates required to ensure the nontriviality of the absolutely continuous spectrum. Moreover, it was also shown that in such a setting there are no eigenvalues in the spectrum. In the case of nonnegative sectional curvature, explicit convex functions have been used to show that there are no eigenfunctions under the additional assumptions on the codimension of the soul of the manifold being $1$ or having a quadratic decay of radial curvatures when the soul is just a point [14, Theorems 4.2, 4.3].
One of the key results of this article is the construction of a $C^{2}$-strictly convex function with bounded gradient on complete simply connected Riemannian manifolds without focal points having negative Ricci curvature (Theorem 3.1). A distinguishing feature here is that we do not make any assumptions on the sign of the sectional curvature. It is well known [18] that there are both compact and noncompact manifolds without focal points whose sectional curvatures change sign. Since the Ricci curvature is an average of sectional curvatures, the sign of the Ricci curvature is a weaker notion when compared to the sign of the sectional curvature. That this is true especially for negative Ricci curvature can be seen from the remarkable result [23, Theorem A] that any smooth manifold of dimension at least three admits a complete Riemannian metric with negative Ricci curvature. Interesting classes of manifolds satisfying our assumptions include Riemannian symmetric spaces of noncompact type as well as simply connected noncompact nonflat harmonic manifolds without focal points. Harmonic manifolds are Einstein [21] i.e, $Ric_{g}=\kappa g$, and have horospheres, which are level sets of Busemann functions, of constant mean curvature $h\geq 0$. If the Ricci curvature $\kappa>0$, then the harmonic manifold is necessarily a compact rank one symmetric space and if $\kappa=0$, then it is necessarily Euclidean. All other simply connected harmonic manifolds have negative Ricci curvature. We remark here that $h>0$ for such manifolds, since a harmonic manifold is Euclidean when $h=0$ [21, §5, Theorem 19].
@@
For the construction of the strictly convex function we make use of Jacobi tensors introduced in [13] and its relation to the Hessian of Busemann functions for manifolds of continuous asymptote. In particular, for manifolds without focal points these Hessians can be seen to be nonnegative. So each Busemann function is convex but not strictly convex since its Hessian vanishes in the direction of the gradient. Motivated by the work done in the context of symmetric spaces of noncompact type in [10], we then average the Busemann functions over the boundary of the geodesic compactification for manifolds without focal points. Here we use [16] for the geometry at infinity of such manifolds. Under the assumption of negative Ricci curvature, a strictly convex $C^{2}$-function with bounded gradient is then obtained. As an immediate consequence, we have that the manifolds under consideration admit smooth strictly convex functions which also has bounded gradient. This follows from the approximation result of Green and Wu [17, Theorem 2(a)]. It is known [26] that the smooth approximation result fails for functions which are merely convex.
Our next application is in the context of manifolds which are, in addition, Kähler. A complex manifold with a complex structure $J$ and a compatible Riemannian metric $g$ is called Kähler if the associated alternating two-form defined by $\omega(X,Y):=g(JX,Y)$ is symplectic. Given any two of these compatible structures the third one can be recovered. ${\mathbb{C}}^{n}$ with the standard metric, bounded domains in ${\mathbb{C}}^{n}$ with the Bergmann metric, Hermitian symmetric spaces with the invariant metric, complex projective spaces with the Fubini-Study metric are some illustrative examples of Kähler manifolds. $C^{2}$-strictly convex functions on Kähler manifolds can be seen to be strictly plurisubharmonic functions. In our context, therefore, we get a smooth strictly plurisubharmonic function with bounded gradient. From this we derive a lower bound for the growth of the volume of metric balls on such manifolds following [24, Lemma 4.2 (i)]. The growth of the volume of the metric balls has been usually studied with lower bounds on Ricci curvature and upper bounds on sectional curvatures [7, §III.4].
As principal applications of Theorem 3.1, we describe the spectrum of the Laplacian for certain classes of manifolds having horospheres of constant mean curvature $h$. Complete simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$ are known as asymptotically harmonic manifolds [21]. On manifolds as in Theorem 3.1, which are in addition asymptotically harmonic, we observe that the strictly convex function that we construct also has bounded Laplacian and proceed to show in Theorem 4.3 that the spectrum is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$. The condition on the bilaplacian required as in [28, Theorem 2] follows automatically since the horospheres have constant mean curvature. In fact, along the way we prove that equality holds in Cheeger’s inequality [6, §4.3, Theorem 3] for asymptotically harmonic manifolds and that Cheeger’s constant is equal to $h$ in Proposition 4.1. Indeed, the proof shows that it is sufficient for horospheres to have constant mean curvature in one direction. Further, we prove in Theorem 4.7 that the spectrum is purely absolutely continuous for simply connected noncompact nonflat harmonic manifolds without focal points and is equal to $[h^{2}/4,\infty)$. These are harmonic manifolds of purely exponential volume growth [21, Theorems 32, 45]. Rank one symmetric spaces of noncompact type and other Damek-Ricci spaces are some examples of this class of manifolds. Since harmonic manifolds are asymptotically harmonic, that the spectrum is purely absolutely continuous and that it is contained in $[h^{2}/4,\infty)$ follow from Theorem 4.3. To prove that the spectrum is equal to $[h^{2}/4,\infty)$, we make use of the Fourier transform developed in this context in [3]. More precisely, using the unitarity of the spherical Fourier transform [3, Theorem 4.7], we can show in the usual way that the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$ is unitarily equivalent to the action of the Laplacian on radial functions to conclude that $[h^{2}/4,\infty)$ is contained in the spectrum.
Higher rank symmetric spaces of noncompact type are not asymptotically harmonic since the horospheres along each direction have constant mean curvature which depend on that direction. Nonetheless, we are able to obtain (Theorem 4.8) that the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\left\lVert\rho^{2}\right\rVert,\infty)$ by making use of the explicit expression for the mean curvature of horospheres and the strictly convex function that we construct in Theorem 3.1. To show that the spectrum is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$, we use the unitarity of just the spherical Fourier transform as in the case of harmonic manifolds. Note that in contrast to the case of harmonic manifolds where the unitarity of the Fourier transform is not known, it is classical that the Fourier transform in the case of symmetric spaces is unitary. As a consequence, the spectrum can also be completely determined for symmetric spaces via the unitary equivalence between the Laplacian and the multiplication operator $M_{(\left\lVert\lambda\right\rVert^{2}+\left\lVert\rho\right\rVert^{2})}$ given by the Fourier transform. It is worth remarking that in showing that the spectrum of the Laplacian on symmetric spaces of noncompact type is purely absolutely continuous, essential use is made only of the fact that they are manifolds without focal points and have negative Ricci curvature.
@@ 2. Preliminaries
Let $M$ be a complete Riemannian manifold and let $\gamma$ be a geodesic in $M$. A vector field $Y$ on $\gamma$ is called a Jacobi field if it satisfies the Jacobi equation
$Y^{{}^{\prime\prime}}+R(\gamma^{{}^{\prime}},Y)\gamma^{{}^{\prime}}=0.$
Here $\gamma^{{}^{\prime}}$ is the tangent vector field of $\gamma$, $R$ is the Riemann curvature tensor on $M$ and $Y^{{}^{\prime}}$ denotes the covariant derivative of $Y$ with respect to $\gamma^{{}^{\prime}}$ along $\gamma$. Two points $p$ and $q$ on $\gamma$ are said to be conjugate along $\gamma$ if there exists a nonzero Jacobi field which vanishes both at $p$ and at $q$. $M$ is called a manifold without conjugate points if for every geodesic $\gamma$ in $M$, no two points on $\gamma$ are conjugate along $\gamma$ or equivalently, if the exponential map is non-singular at every point in $M$. In fact, $M$ has no conjugate points if and only if $\exp_{p}:T_{p}M\to M$ is a covering map for each $p$ in $M$. So if $M$ is also simply connected, then $\exp_{p}$ is a diffeomorphism for all $p\in M$ and hence $M$ is necessarily noncompact. In this case, any two distinct points of $M$ are joined by a unique geodesic. A general reference for the basic Riemannian geometry used in this article is [7].
We now give a brief overview of Jacobi tensors and refer to [13] for a full account. Let $M$ be a complete simply connected Riemannian manifold without conjugate points and let ($SM$) $TM$ denote its (unit) tangent bundle. For $v\in SM$, let $\gamma_{v}$ denote the unit speed geodesic determined by $v$. A normal $(1,1)$-tensor $Y_{v}$ along the geodesic $\gamma_{v}$ is a bundle endomorphism of the normal bundle $N\gamma_{v}$ over $\gamma_{v}$ given by
@@
A Jacobi tensor $Y_{v}$ is a normal $(1,1)$-tensor field along a geodesic $\gamma_{v}$ with transversal derivative (i.e. there exists some $t$ such that $ker~{}Y_{v}(t)\cap ker~{}Y_{v}^{{}^{\prime}}(t)=0$) satisfying the Jacobi equation
(2.1) $Y_{v}^{{}^{\prime\prime}}(t)+R(t)Y_{v}(t)=0$
@@
$R(t)w:=R(w,\gamma_{v}^{{}^{\prime}}(t))\gamma_{v}^{{}^{\prime}}(t)$
and $Y_{v}^{{}^{\prime}}$ is the covariant derivative of the tensor $Y_{v}$ along $\gamma_{v}$ with respect to the tangent vector field $\gamma_{v}^{{}^{\prime}}$. For each $s>0$ and $v\in SM$, let $D_{s,v}$ be a solution to (2.1) with the boundary conditions $D_{s,v}(0)=\mathrm{Id}$ and $D_{s,v}(s)=0$. Then $D_{v}:=\lim\limits_{s\to\infty}D_{s,v}$ exists and is referred to as the stable Jacobi tensor $D_{v}$ satisfying $D_{v}(0)=\mathrm{Id}$. A complete simply connected Riemannian manifold without conjugate points is said to be manifold of continuous asymptote if $v\to D_{v}^{{}^{\prime}}(0)$ is a continuous section of the vector bundle $\{(v,\phi):v\in SM,\phi\in\mathrm{End}(N(\gamma_{v})(0))\}$ over $SM$.
Given $t\in[0,\infty)$ and $v\in SM$, let $b_{t,v}:M\to\mathbb{R}$ denote the function defined by
@@
is defined everywhere on $M$. The function $b_{v}$ is called the Busemann function associated to $v\in SM$. Recall that the level sets of Busemann functions are called horospheres. Horospheres are closed and hence are complete hypersurfaces with the induced metric. Recall that for Hadamard manifolds the Busemann functions are $C^{2}$ [12, §1.6.6]. Eschenburg [13] was able to relax the curvature restriction by making use of convergence properties of Jacobi tensors, and generalized the $C^{2}$-regularity of horospheres to manifolds of continuous asymptote. For such manifolds, [22, §11] gives a compactification known as the Busemann compactification. Let $o\in M$ be an arbitrary but fixed point. Then the Busemann boundary $\partial^{o}M$ is homeomorphic to $S_{p}M$ for any $p\in M$ [22, Proposition 11.4 (2)].
We see that for each $v\in SM$, the Hessian of the Busemann function $b_{v}$ and the stable Jacobi tensor $D_{v}$ are related as follows. The orthogonal projection $x^{\perp}(p,v)$ of $x\in T_{p}M$ onto $(\nabla b_{v}(p))^{\perp}$ will be denoted by $x^{\perp}$ for notational convenience.
@@
For $p\in M$ and $x\in S_{p}M$, we have ${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp})$.

### v2

# Convexity on manifolds without focal points and applications

@@ Abstract.
In this article, we study strict convexity on simply connected Riemannian manifolds without focal points, a class of manifolds which contains not only Hadamard manifolds but also examples of manifolds whose sectional curvatures change sign. Using the geometrically defined convex functions on such manifolds, namely the distance function and the Busemann function, strictly convex functions with various desired additional properties are constructed. From the former, we derive interesting consequences such as the continuity of the isoperimetric profile function without conditions on the sectional curvature and, if the manifolds are also Kähler, Steinness as well as a lower bound on the volume growth of metric balls. Our primary applications concern the spectrum of the Laplacian. We prove that the absolutely continuous part of the spectrum contains a certain infinite interval assuming only the existence of a point with respect to which the radial curvatures are nonpositive. This is a substantial generalisation of the corresponding result for Hadamard manifolds. In the latter case of Busemann functions, we use the geometry at infinity to give a new construction of a strictly convex function assuming that the Ricci curvature is negative. We then apply this to show that the spectrum is purely absolutely continuous on a class of manifolds for which all the horospheres have the same constant mean curvature. In particular, for a well-studied subclass for which the condition on Ricci curvature is automatically satisfied, we completely determine the spectrum. This construction also enables us to recover the classical result on the spectrum in the case of symmetric spaces of noncompact type and higher rank even though the mean curvatures of horospheres there are not constant.
@@ 1. Introduction
Convex functions have long played a fundamental role in Riemannian geometry. By the convexity of a continuous real-valued function on a Riemannian manifold we mean that it is convex in the usual sense along any geodesic. The mere existence or nonexistence of convex functions have geometric and topological implications. Nonconstant convex functions on complete Riemannian manifolds exist only on those with infinite volume. In particular, they don’t exist on compact Riemannian manifolds. There are profound structural results arising from the interaction of curvature and convexity. Under appropriate curvature conditions there are natural geometrically defined convex functions on a Riemannian manifold. On simply connected Riemannian manifolds of nonpositive sectional curvature, called Hadamard manifolds, the square of the distance function from a fixed but arbitrary point and the Busemann functions are examples of convex functions which play a crucial role in the geometry of these manifolds. While the former is a smooth strictly convex function, the latter is merely a $C^{2}$-convex function [Ebe96, §1.6.6]. In the nonnegative sectional curvature situation for instance, a key idea in the fundamental work [CG72] is the construction of suitable continuous convex functions related to Busemann functions and having appropriate additional properties.
@@
A function is (strictly) convex if its Hessian is positive (definite) semidefinite everywhere. If a Riemannian manifold admits a strictly convex function, then it has no compact immersed minimal submanifolds. In particular, it has no closed geodesics. When the strictly convex function has a minimum, such a manifold also allows for a generalisation of Cartan’s theorem [Yam82]: every compact subgroup of the isometry group has a common fixed point. The existence of strict convex functions also has consequences for the $\mathrm{L^{2}}$-spectrum of the Laplacian. The study of the spectrum of the Laplacian on complete Riemannian manifolds has a long, rich history and there are very many results, often with various additional constraints placed on the structure of the manifold and on its curvature. For instance, in the case of manifolds of nonpositive curvature the essential spectrum was shown to be an infinite interval depending on the dimension and sectional curvatures at infinity in [Don81, Theorem 6.3] using the distance function. One can then ask interesting questions about the finer structure of the spectrum e.g., whether there are eigenvalues in the essential spectrum. For a succinct recent overview of some of these aspects, we refer to [Don10]. We now content ourselves by mentioning a few results which use convexity in an essential manner to deduce refined information on the nature of the spectrum and to put our work in context, with no pretensions to completeness. In the case of nonnegative sectional curvature, explicit convex functions have also been used to show that there are no $\mathrm{L}^{2}$-eigenfunctions under the additional assumptions on the codimension of the soul of the manifold being $1$ or having a quadratic decay of radial curvatures when the soul is just a point [EF92, Theorems 4.2, 4.3]. On Hadamard manifolds, [Xav88, Theorem 1] exhibited a strictly convex function related to the square of the distance function satisfying the estimates required to ensure the nontriviality of the absolutely continuous spectrum. However, this is not sufficient to exclude the existence either of eigenvalues or of the singularly continuous spectrum.
In this article, we consider convex functions on complete Riemannian manifolds $M$ whose Jacobi fields satisfy certain growth properties, while making no assumptions on sectional curvatures. A simply connected Riemannian manifold is said to have no focal points if the norm of any nontrivial Jacobi field which vanishes at the initial point of a geodesic ray is strictly increasing. Since the exponential map at any point on such a manifold is a diffeomorphism, i.e. every point is a pole, they are necessarily noncompact. This class contains all Hadamard manifolds; however, it is well known [Gul75] that there are both compact and noncompact manifolds without focal points whose sectional curvatures change sign. While several aspects of the geometry of these manifolds have been well-studied, from understanding the dynamics of the geodesic flow [Ebe72], [O’S74], [Esc77]) to recent work on X-ray transforms on such manifolds with strictly convex boundary [PSUZ19] and the solution to the Dirichlet problem at infinity [LZ25], some aspects such as the spectrum of the Laplacian have not been analysed as much. We study some of these using the convexity of two kinds of geometrically defined functions on them. One of them is the square of the distance function from an arbitrary but fixed point. This is a smooth strictly convex function [Ebe72, Lemma 1] whose sub-level sets are compact. It is suitably modified to ensure a bounded gradient, an aspect that is pivotal in many applications. The other is the Busemann function associated to a unit vector in the tangent bundle of $M$. These are $C^{2}$ convex functions [Esc77, Theorem 2] but are not strictly convex. We remedy the situation by giving a new construction of a strictly convex function from Busemann functions. This leads to some new spectral theoretic results. A careful analysis of Jacobi fields and Jacobi tensors is central to these considerations. We now proceed to discuss our results in some detail.
Our first application using the strict convexity of the suitably modified distance function concerns the isoperimetric profile function $\mathcal{I}_{M}$ defined on the manifold $M$. This function maps every positive volume to the infimum of the perimeters of sets in $M$ having that volume. The isoperimetric profile is shown [Rit17, Theorem 3.1] to be continuous and nondecreasing for complete manifolds which possess smooth strictly convex Lipschitz exhaustion functions and is therefore differentiable almost everywhere on such manifolds. Examples are discussed under sectional curvature conditions there, namely, Hadamard manifolds and manifolds with strictly positive sectional curvatures. In contrast, we observe in Proposition 3.2 that for simply connected Riemannian manifolds without focal points the map $\mathcal{I}_{M}$ is a nondecreasing continuous function, thus giving a class of examples without constraints on sectional curvatures.
Strictly convex functions on Kähler manifolds can be seen to be strictly plurisubharmonic functions. So, manifolds without focal points which are also Kähler, admit strictly plurisubharmonic functions obtained, for instance, via the square of the distance function. From this, we observe in Proposition 3.4 that such manifolds are Stein using Grauert’s proof of Levi’s problem [GW77, §4]. By composing with a suitable strictly convex function on $\mathbb{R}$, we get a strictly convex exhaustion function with bounded gradient. This is applied in Proposition 3.5 to obtain a lower bound for the growth of the volume of metric balls following [NT03, Lemma 4.2 (i)]. The growth of the volume of the metric balls has been usually studied with lower bounds on Ricci curvature and upper bounds on sectional curvatures. It is an open problem whether a simply connected harmonic manifold which admits a Kähler structure is a symmetric space. Indeed in the Damek-Ricci setting this is known [BTV95, §4.1.12]. Therefore, at least in certain special cases such as harmonic manifolds without focal points, this volume growth is in line with the above mentioned rigidity by Bishop-Günther’s comparison theorem [Cha06, Theorem III.4.2].
We now proceed to describe our first spectral-theoretic result and note that the smooth strictly convex function with bounded gradient discussed in the previous paragraph is used here as well. A key idea is due to [Xav88] showing that on complete Riemannian manifolds supporting a sufficiently regular strictly convex function, Kato’s theory of $H$-smooth operators can be applied to understand the spectrum of the Laplacian. For a detailed discussion on $H$-smoothness we refer to the comprehensive survey [Sim19, §4]. In the case of Hadamard manifolds, this was done in [Xav88, Theorem 1] under suitable bounds on the Laplacian and its derivatives to conclude that the absolutely continuous part of the spectrum was nontrivial. In fact, it contains an interval that depends explicitly on these bounds. In the present work, we extend the implementation of $H$-smoothing to simply connected Riemannian manifolds without focal points in Theorem 3.8. Assuming only the nonpositivity of radial curvatures with respect to a point, a condition that is markedly weaker than global nonpositivity of sectional curvature, and bounds on Laplacian and the bilaplacian, we show that the absolutely continuous spectrum is an infinite interval which depends on the bounds. This is carried out by applying Lemma 3.6 where we obtain the required quantitative estimates on the Hessian of distance function in the usual fashion, while using a function that is defined when the manifold necessarily has no focal points. For context, we mention some of the work that derive information about the spectrum from the behaviour of the radial curvatures. Absence of point spectrum has been shown for manifolds with a pole satisfying quadratic decay of radial curvatures [Don10, Theorem 1.5]. Under linear decay of radial curvatures it was proved in [Kum10] that there are no eigenvalues embedded in the essential spectrum. If the radial curvature is assumed to approach a strictly negative constant near infinity, then [JL20] investigated the nature of the singular spectrum.
In view of Theorem 3.8, we can ask for more refined properties of the spectrum such as identifying whether the point spectrum or the singularly continuous spectrum are nontrivial. A sufficient condition for the absence of point spectrum and the singularly continuous spectrum is given by the existence of a strictly convex function whose bilaplacian is nonpositive [Xav88, Corollary to Theorem 2]. As noted there, this seemingly unnatural additional condition is, in fact, necessary. Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed. This sufficiency condition seems to us not to be approachable by the use of distance function.
@@
To overcome this, we now turn our attention on the other natural geometrically defined convex function on simply connected Riemmanian manifolds without focal points, the Busemann function. Note that the Busemann function is convex but not strictly convex due to the vanishing of Hessian in the direction of its gradient. We use the geometry at infinity to construct a strictly convex function in Theorem 4.1 by averaging over all Busemann functions. This was motivated by the idea outlined in the context of symmetric spaces of noncompact type in [DX06, p-657], and is possibly of independent interest. For this, we make use of Jacobi tensors introduced in [Esc77] and its relation to the Hessian of Busemann functions for manifolds of continuous asymptote. In particular, for manifolds without focal points these Hessians can be seen to be positive semidefinite. The Hessian of the convex function which we construct by averaging over Buseman functions becomes positive definite under the assumption of negative Ricci curvature, yielding Theorem 4.1.
We now examine the hypotheses of Theorem 4.1. In this class of manifolds, having nonnegative Ricci curvature at even one point, forces the manifold to be flat [MZ00, Corollary 1]. Being an average of sectional curvatures, the sign of the Ricci curvature is a milder hypothesis. That this is true especially for negative Ricci curvature can be seen from the remarkable result [Loh94, Theorem A] that any smooth manifold of dimension at least three admits a complete Riemannian metric with negative Ricci curvature. Well-known classes of manifolds satisfying our assumptions include Riemannian symmetric spaces of noncompact type as well as simply connected noncompact (nonflat) harmonic manifolds without focal points. Harmonic manifolds are Einstein [Kni16] i.e, $Ric_{g}=\kappa g$, and have horospheres, which are level sets of Busemann functions, of constant mean curvature $h\geq 0$. If the Ricci curvature $\kappa>0$, then the harmonic manifold is necessarily a compact rank one symmetric space and if $\kappa=0$, then it is necessarily Euclidean. All other simply connected harmonic manifolds have negative Ricci curvature. Rank one symmetric spaces of noncompact type and other Damek-Ricci spaces are examples of this class of manifolds. We remark here that $h>0$ for such manifolds, since a harmonic manifold is Euclidean when $h=0$ [Kni16, §5, Theorem 19]. Harmonic manifolds without focal points are of purely exponential volume growth, and these have been well studied in the recent past (e.g.[Kni16, BKP21, Bra24]).
As principal applications of Theorem 4.1, we fully describe the nature of the spectrum $\sigma(\Delta)$ of the Laplacian $\Delta$ for interesting classes of manifolds having horospheres of constant mean curvature $h$. Simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$ are known as asymptotically harmonic manifolds [Kni16]. For all asymptotically harmonic manifolds, we show that equality holds in Cheeger’s inequality and that the Cheeger’s constant is equal to $h$ (see [KP15, §5] for related results). We mention in this context that for asymptotically harmonic Hadamard manifolds, the essential spectrum has been determined in [CS14, Theorem 4.4]. But it did not address finer questions such as existence of embedded eigenvalues or the singularly continuous spectrum. It was proved in [BMP23] that for asymptotically harmonic Hadamard manifolds with strictly negative sectional curvature (and so necessarily without focal points) the spectrum is absolutely continuous. On simply connected manifolds without focal points and negative Ricci curvature, as in Theorem 4.1, which are asymptotically harmonic, we show in Theorem 4.3 that the spectrum is purely absolutely continuous and is contained in $h^{2}/4,\infty)$.
Now since harmonic manifolds are asymptotically harmonic, the above result also holds in this case (Corollary 4.7). We give an elegant proof of the equality $\sigma(\Delta)=[h^{2}/4,\infty)$, making use of the Fourier transform developed recently in the context of harmonic manifolds without focal points in [BKP21]. It is not known whether this Fourier transform is unitary on the space of all $\mathrm{L}^{2}$-functions. However, the Fourier transform is unitary on radial $\mathrm{L}^{2}$-functions [BKP21, Theorem 4.7] with the Plancherel density given by the $\mathrm{c}$-function, and intertwines the radial Laplacian and the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$. Using certain estimates on the $\mathrm{c}$-function, we show that the essential range of the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$ is equal to $h^{2}/4,\infty)$, enabling us to conclude the desired equality for the spectrum.
Although symmetric spaces of higher rank are not asymptotically harmonic, our geometric methods allow us to recover the absolute continuity of the spectrum. It is, of course, classical that the unitarity of the Fourier transform in the case of symmetric spaces of noncompact type completely determines the spectrum of the Laplacian to be absolutely continuous and equal to $\left\lVert\rho^{2}\right\rVert,\infty)$.
@@ 2. Preliminaries
In this section, we give a brief summary of the basic geometric setup needed in the sequel. Let $M$ be a complete Riemannian manifold and let $\gamma$ be a geodesic in $M$. A vector field $Y$ along $\gamma$ is called a Jacobi field if it satisfies the Jacobi equation
(1) $Y^{{}^{\prime\prime}}+R(\gamma^{{}^{\prime}},Y)\gamma^{{}^{\prime}}=0.$
Here $\gamma^{{}^{\prime}}$ is the tangent vector field of $\gamma$, $R$ is the Riemann curvature tensor on $M$ and $Y^{{}^{\prime}}$ denotes the covariant derivative of $Y$ with respect to $\gamma^{{}^{\prime}}$ along $\gamma$. The prototypical example of a Jacobi field is the variation vector field of a geodesic variation. Two points on $\gamma$ are said to be conjugate along $\gamma$ if there exists a nonzero Jacobi field which vanishes at both the points. $M$ is called a manifold without conjugate points if for every geodesic $\gamma$ in $M$, no two points on $\gamma$ are conjugate along $\gamma$ or equivalently, if the exponential map is non-singular at every point in $M$. In fact, $M$ has no conjugate points if and only if $\exp_{p}:T_{p}M\to M$ is a covering map for each $p$ in $M$. So if $M$ is also simply connected, then $\exp_{p}$ is a diffeomorphism for all $p\in M$ [O’S74, Theorem 1] and hence $M$ is necessarily noncompact. In this case, any two distinct points of $M$ are joined by a unique geodesic.
A simply connected Riemannian manifold $M$ is said to have no focal points if for every nontrivial Jacobi field $Y$ satisfying $Y(0)=0$, the map $t\mapsto\left\lVert Y(t)\right\rVert$ is strictly increasing for $t>0$ [O’S74, Proposition 4]. In particular, any non-trivial Jacobi field cannot vanish more than once and hence every manifold without focal points is also a manifold without conjugate points. Note that all Hadamard manifolds are manifolds without focal points.
For some applications, it is not just convenient but even important to consider the full vector space of normal Jacobi fields along a unit speed geodesic together rather than consider them one at a time. A Jacobi tensor is the tool which does this. We now give a brief overview of Jacobi tensors and refer to [Esc77] for a full account. Let $M$ be a simply connected Riemannian manifold without conjugate points and let ($SM$) TMTM denote its (unit) tangent bundle. For $v\in SM$, let $\gamma_{v}$ denote the unit speed geodesic determined by $v$. A normal $(1,1)$ - tensor $Y_{v}$ along the geodesic $\gamma_{v}$ is a bundle endomorphism of the normal bundle $N\gamma_{v}$ over $\gamma_{v}$ given by
@@
A Jacobi tensor $Y_{v}$ is a normal $(1,1)$ - tensor field along a geodesic $\gamma_{v}$ with transversal derivative (i.e. there exists some $t$ such that $ker\penalty 10000\ Y_{v}(t)\cap ker\penalty 10000\ Y_{v}^{{}^{\prime}}(t)=0$) satisfying the Jacobi equation
(2) $Y_{v}^{{}^{\prime\prime}}(t)+R(t)Y_{v}(t)=0$
@@
$R(t)w:=R(\gamma_{v}^{{}^{\prime}}(t),w)\gamma_{v}^{{}^{\prime}}(t)$
and $Y_{v}^{{}^{\prime}}$ is the covariant derivative of the tensor $Y_{v}$ along $\gamma_{v}$ with respect to the tangent vector field $\gamma_{v}^{{}^{\prime}}$. For each $s>0$ and $v\in SM$, let $D_{s,v}$ be a solution to the Jacobi equation (2) with the boundary conditions $D_{s,v}(0)=\mathrm{Id}$ and $D_{s,v}(s)=0$. Then $D_{v}:=\lim\limits_{s\to\infty}D_{s,v}$ exists and is referred to as the stable Jacobi tensor $D_{v}$ satisfying $D_{v}(0)=\mathrm{Id}$. A simply connected Riemannian manifold without conjugate points is said to be a manifold of continuous asymptote if $v\to D_{v}^{{}^{\prime}}(0)$ is a continuous section of the vector bundle $\{(v,\phi):v\in SM,\phi\in\mathrm{End}(N(\gamma_{v})(0))\}$ over $SM$. As stated explicitly in [Got79, p-596] simply connected manifolds without focal points fall in this class. For such manifolds, [KP13, §11] gives a compactification known as the Busemann compactification. Let $o\in M$ be an arbitrary but fixed point. Then the Busemann boundary $\partial^{o}M$ is homeomorphic to $S_{p}M$ for any $p\in M$ [KP13, Proposition 11.4 (2)].
Convex functions being the theme of this article, the two natural geometrically defined functions which under suitable constraints on curvature are convex are the distance function and the Busemann function defined below. Given $t\in[0,\infty)$ and $v\in SM$, let $b_{t,v}:M\to\mathbb{R}$ denote the function defined by
@@
is defined everywhere on $M$. The function $b_{v}$ is called the Busemann function associated to $v\in SM$. The level sets of Busemann functions are called horospheres. Horospheres are closed and hence are complete hypersurfaces with the induced metric. It is classical that [Ebe96, §1.6.6] for Hadamard manifolds the Busemann functions are $C^{2}$ convex functions. Eschenburg [Esc77] was able to relax the curvature restriction by making use of convergence properties of Jacobi tensors, and generalized the $C^{2}$ - regularity of horospheres to manifolds of continuous asymptote. In order to obtain this, the following relation between the Hessian of the Busemann function $b_{v}$ and the stable Jacobi tensor $D_{v}$ for each $v\in SM$ is used. We first make a small observation. The orthogonal projection $x^{\perp}(p,v)$ of $x\in T_{p}M$ onto $(\nabla b_{v}(p))^{\perp}$ will be denoted by $x^{\perp}$ for notational convenience.
@@
For $p\in M$ and $x\in S_{p}M$, we have $\mathrm{Hess}\penalty 10000\ b_{v}(p)(x,x)=\mathrm{Hess}\penalty 10000\ b_{v}(p)(x^{\perp},x^{\perp})$.